# A nonlinear real time capable motion cueing algorithm based on deep reinforcement learning

### v1

@@ III-B Algorithms for Validation
In this chapter, the state-of-the-art algorithms, including CW, linear MPC (LMPC), and nonlinear MPC (NMPC) that are being used for validation are presented.
@@
For the LMPC based MCA used here plant models formulated in Cartesian space are used which consist of single integrators for the rotational degrees of freedom and double integrators for the translational degrees of freedom. The plant inputs are velocity for rotational degrees of freedom and acceleration for translational degrees of freedom. The underlying optimization problem of the controllers consists of minimizing the sensation error with respect to a given reference signal as well as minimizing the Cartesian position, velocity and acceleration of the robot without violating any prescribed limits. These control objectives are implemented as a quadratic cost function with linear constraints, a formulation that is well suited for real-time optimization. At every control iteration the output of the optimizer is a set of acceleration and velocity profiles to follow. These profiles are integrated into position commands. The implementation of this algorithm is done in MATLAB using the optimal control toolbox.
@@ III-C Results of Validation
To assess the performance of the DRL MCA, a comprehensive comparison with the CW, LMPC and NMPC MCAs is conducted. For this purpose, 15 different validation files, each spanning 51 seconds, are recorded. These files include a range of driving scenarios, such as normal driving, harsh braking and acceleration, slalom driving, and continuous circular driving. The 15 validation files serve as input for all four algorithms. To ensure a realistic comparison, the trajectories generated by each MCA are applied to a model of the robotic MSP. This model limits the joint positions, velocities, and accelerations to reproduce a realistic behavior of the MSP and is similar to the model used during the training of the DRL MCA and in the optimization of the NMPC MCA. In case of strong violations of the limitations in joint positions, velocities or acceleration this post processing reduces the quality of the motion sensation severely as a strong deviation in suggested and physically possible trajectory is induced. Therefore, CW and LMPC in combination with the necessary post processing, achieve the best possible result with a conservative tuning, as only minor changes to the suggested trajectory become necessary.
@@
All objective values are presented in Table II. To gain a more comprehensive understanding of the results, a Wilcoxon Signed-Rank test was performed. This statistical test is a non-parametric alternative to the paired t-test and is used to compare two paired samples. The test was conducted to compare the performance of the DRL MCA against each of the three other algorithms, allowing for an assessment of the DRL MCA’s performance. For all three comparisons, the null hypothesis states that there is no significant difference of the objective values between the two algorithms being compared. The significance level ($\alpha$) for rejecting the null hypothesis was set to $0.05$. Both comparisons, DRL MCA to CW MCA and DLR MCA to LMPC MCA, resulted in a p-value of $6.104\cdot 10^{-5}$, indicating clearly that the null hypothesis can be rejected, and a significant difference in performance exists between DRL and the two algorithms. As shown in Table II, the DRL MCA achieved better objective values than CW MCA and LMPC MCA across all 15 trajectories. The third Wilcoxon Signed-Rank test compared the DRL MCA to the NMPC MCA and resulted in a p-value of $0.561$. This indicates that the null hypothesis cannot be rejected for this test, and no significant difference in performance was found between these two algorithms. The three conducted Wilcoxon Signed-Rank tests lead to the conclusion that CW MCA and LMPC MCA show significant worse result than DRL MCA. NMPC shows similar performances based on the analyzed metric.
@@
Figure 5: Specific force and angular velocity for the validation trajectory 3. Presented are the values of the reference and the output of the NMPC, LMPC, CW and DRL MCAs.
@@
Next, a more detailed assessment of one single trajectory is conducted. Trajectory 3 is chosen here as it provides comparably similar objective values for NMPC and DRL MCA. Figure 5 provides a qualitative assessment by comparing the specific forces and angular velocities produced by the four algorithms to the reference trajectory for validation trajectory 3. Combined with Table III and Table IV, presenting the RMSEs and PCCs for this trajectory across all four algorithms, a more detailed evaluation is possible. Comparing the performance of DRL MCA and NMPC MCA on the angular velocities, $\boldsymbol{\omega}_{x}$ and $\boldsymbol{\omega}_{y}$, against the performance on $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$ it becomes clear that the amount of tilt coordination differs strongly. This impression gets confirmed by looking at the RMSE and the PCC: NMPC MCA achieves a better match in $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$ by performing more tilt coordination by rotational movement of the simulator cabin, which leads to stronger deviations in $\boldsymbol{\omega}_{x}$ and $\boldsymbol{\omega}_{y}$. Conversely, DLR MCA performs less tilt coordination, so the angular velocities show a better match but the specific forces are followed less closely. This is a typical trade off in the design of MCAs and especially for driving simulations, that come with a more dynamic acceleration profile than flight simulation, the preferred rate of tilt coordination is highly subjective and an optimal value is not determined [43, 42]. The performance of LMPC and CW on the trajectory is unsatisfactory as both algorithms are only able to follow the reference with deviation in the phase, particularly noticeable for $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$. This can be assigned to a conservative tuning that becomes necessary as the path planning is confined to the Cartesian workspace without consideration of the limitations of the MSP’s joints.
@@
Figure 6 illustrates the position and orientation of the MSP’s end effector for trajectory 3 across all four algorithms. It is evident that NMPC utilizes tilt coordination the most, as indicated by the significant changes in orientation. Overall, NMPC makes greater use of the workspace compared to the other three algorithms. This highlights that the DRL MCA achieves comparable results for specific force and angular velocity while utilizing the resources of the MSP more efficiently.
@@
Figure 6: Position and orientation of the end effector for the validation trajectory 3. Presented are the values of NMPC, LMPC, CW and DRL MCAs.
@@
Algorithms CW LMPC NMPC DRL
Average Time [ms] 0.300 0.012 1546.5 0.965
Looking at the computational load, as presented in Table V, the strength of the DRL MCA becomes clear. The average computational time per time step of $12$ ms is $0.965$ ms ($\sigma=0.501$ ms), making the algorithm clearly real time capable. NMPC MCA, being the only other algorithm which considers the model of the MSP and therefore is generating a feasible trajectory for the robotic simulator, reaches an average value of $1546.5$ ms ($\sigma=460.7$ ms) per $12$ ms time step. This high computational load makes it not real time capable and not applicable to the real system. With $0.300$ ms and $0.012$ ms LMPC and CW are also able to meet the real-time requirements, due to their comparatively simple structure, which does not take into account the nonlinear dynamics of the MSP. However, as shown in the previous discussion, they are not able to achieve compatible performance in this setup.

### v2

@@ III-B Algorithms for Validation
In this chapter, the state-of-the-art algorithms, including CW and nonlinear MPC (NMPC) that are being used for validation are presented.
@@ III-C Results of Validation
To assess the performance of the DRL MCA, a comprehensive comparison with the CW and NMPC MCAs is conducted. For this purpose, 15 different validation files, each spanning 51 seconds, are recorded. These files include a range of driving scenarios, such as normal driving, harsh braking and acceleration, slalom driving, and continuous circular driving. The 15 validation files serve as input for all four algorithms. To ensure a realistic comparison, the trajectories generated by each MCA are applied to a model of the robotic MSP. This model limits the joint positions, velocities, and accelerations to reproduce a realistic behavior of the MSP and is similar to the model used during the training of the DRL MCA and in the optimization of the NMPC MCA. In case of strong violations of the limitations in joint positions, velocities or acceleration this post processing reduces the quality of the motion sensation severely as a strong deviation in suggested and physically possible trajectory is induced. Therefore, CW in combination with the necessary post processing, achieves the best possible result with a conservative tuning, as only minor changes to the suggested trajectory become necessary.
@@
All objective values are presented in Table II. To gain a more comprehensive understanding of the results, a Wilcoxon Signed-Rank test was performed. This statistical test is a non-parametric alternative to the paired t-test and is used to compare two paired samples. The test was conducted to compare the performance of the DRL MCA against both other algorithms, allowing for an assessment of the DRL MCA’s performance. For both comparisons, the null hypothesis states that there is no significant difference of the objective values between the two algorithms being compared. The significance level ($\alpha$) for rejecting the null hypothesis was set to $0.05$. The comparison DRL MCA to CW MCA resulted in a p-value of $6.104\cdot 10^{-5}$, indicating clearly that the null hypothesis can be rejected, and a significant difference in performance exists between DRL MCA and the washout based algorithm. As shown in Table II, the DRL MCA achieved better objective values than CW MCA across all 15 trajectories. The third Wilcoxon Signed-Rank test compared the DRL MCA to the NMPC MCA and resulted in a p-value of $0.561$. This indicates that the null hypothesis cannot be rejected for this test, and no significant difference in performance was found between these two algorithms. The three conducted Wilcoxon Signed-Rank tests lead to the conclusion that CW MCA and LMPC MCA show significant worse result than DRL MCA. NMPC shows similar performances based on the analyzed metric.
@@
Figure 5: Specific force and angular velocity for the validation trajectory 3. Presented are the values of the reference and the output of the NMPC, CW and DRL MCAs.
@@
Next, a more detailed assessment of one single trajectory is conducted. Trajectory 3 is chosen here as it provides comparably similar objective values for NMPC and DRL MCA. Figure 5 provides a qualitative assessment by comparing the specific forces and angular velocities produced by the four algorithms to the reference trajectory for validation trajectory 3. Combined with Table III and Table IV, presenting the RMSEs and PCCs for this trajectory across all four algorithms, a more detailed evaluation is possible. Comparing the performance of DRL MCA and NMPC MCA on the angular velocities, $\boldsymbol{\omega}_{x}$ and $\boldsymbol{\omega}_{y}$, against the performance on $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$ it becomes clear that the amount of tilt coordination differs strongly. This impression gets confirmed by looking at the RMSE and the PCC: NMPC MCA achieves a better match in $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$ by performing more tilt coordination by rotational movement of the simulator cabin, which leads to stronger deviations in $\boldsymbol{\omega}_{x}$ and $\boldsymbol{\omega}_{y}$. Conversely, DLR MCA performs less tilt coordination, so the angular velocities show a better match but the specific forces are followed less closely. This is a typical trade off in the design of MCAs and especially for driving simulations, that come with a more dynamic acceleration profile than flight simulation, the preferred rate of tilt coordination is highly subjective and an optimal value is not determined [43, 42]. The performance of CW on the trajectory is unsatisfactory as it is only able to follow the reference with deviation in the phase, particularly noticeable for $\boldsymbol{f}_{x}$ and $\boldsymbol{f}_{y}$. This can be assigned to a conservative tuning that becomes necessary as the path planning is confined to the Cartesian workspace without consideration of the limitations of the MSP’s joints but is also an inherent issue of the filter based approach.
@@
Figure 6 illustrates the position and orientation of the MSP’s end effector for trajectory 3 across all three algorithms. It is evident that NMPC utilizes tilt coordination the most, as indicated by the significant changes in orientation. Overall, NMPC makes greater use of the workspace compared to the other two algorithms. This highlights that the DRL MCA achieves comparable results for specific force and angular velocity while utilizing the resources of the MSP more efficiently.
@@
Figure 6: Position and orientation of the end effector for the validation trajectory 3. Presented are the values of NMPC, CW and DRL MCAs.
@@
Algorithms CW NMPC DRL
Average Time [ms] 0.300 1546.5 0.965
Looking at the computational load, as presented in Table V, the strength of the DRL MCA becomes clear. The average computational time per time step of $12$ ms is $0.965$ ms ($\sigma=0.501$ ms), making the algorithm clearly real time capable. NMPC MCA, being the only other algorithm which considers the model of the MSP and therefore is generating a feasible trajectory for the robotic simulator, reaches an average value of $1546.5$ ms ($\sigma=460.7$ ms) per $12$ ms time step. This high computational load makes it not real time capable and not applicable to the real system. With $0.300$ ms CW is also able to meet the real-time requirements, due to its comparatively simple structure, which does not take into account the nonlinear dynamics of the MSP. However, as shown in the previous discussion, it is not able to achieve compatible performance in this setup.